Problem: of SN,=? Warden inverson NHR % SN₂= ? a NH2 Erors [a] 30 ml de [agas oro NH gam 91/6 oll
Answer #1

The SN2 reaction produces, as is written, Walden inversion, so the product of pure SN2 would be:

phpsBQFlF.png

Which has a rotating power of +30° and will be called the (+) enantiomer.

If we had pure SN1, we would obtain a racemic mixture of both enantiomers, since it occurs via an intermediate plane carbocation and posterior attack on any of the two faces of the intermediate. Notice that all the (-) enantiomer must come from this reaction, since the SN2 mechanism yields only the (+) enantiomer.

We can first calculate the fraction of each form present in the mixture using:

(+) 30° - 3(-) -30° = 15°

And:

7(+) + 3(-) = 1

From the first expression, we can rearrange:

X(+) = 0.5 + 3(-

We replace this in the second expression:

0.5+1(-) + I(-) = 1

So we have:

2.0 (-) = 0.5

So, x(-) = 0.25 and x(+) = 0.75.

Since all x(-) must come from an SN1 mechanism, we know we have 0.25 of the (-) enantiomer and 0.25 of the (+) enantiomer from SN1, and the remaining 0.5 of the (+) enantiomer from SN2. This means that the percentages of both mechanisms is 50%, since the proportion of product from both is the same.